# Adaptive finite element interpolated neural networks

### v1

@@ 5.1.2. The effect of preconditioning
We consider the preconditioner $\mathbf{B}_{\rm inv\_lin}$, which is defined as the discrete Laplacian resulting from discretisation with the linearised space $V_{h}$ used both as trial and test FE spaces. Thus, $\mathbf{B}_{\rm inv\_lin}$ is a sparse symmetric positive definite (SPD) matrix with reduced computational cost compared to a regular sparse matrix. Note that we apply $k_{U}$ levels of uniform refinement for the partition associated to $U_{h}$ to obtain the mesh for the linearised test space $V_{h}$. Besides, in [17], we have observed the effectiveness of a GMG preconditioner as an approximation to the inverse of $\mathbf{B}_{\rm inv\_lin}$, suggesting its potential benefits for this problem as well. Nonetheless, for the purpose of this study, we just exactly invert $\mathbf{B}_{\rm inv\_lin}$ (by solving linear systems), as it is sufficient to demonstrate the impact of preconditioning. In these experiments, we cut the number of training iterations by reducing those used in the unpreconditioned experiments by a factor of 4, except for the final adaptation step, where these are divided by a factor of 2. This adjustment accommodates for the increased demand for training on the finest meshes.
@@
In addition to evaluating to what extent preconditioners can expedite convergence, we also investigate the impact of norm choice on the convergence. We train the same NN (i.e., the same parameter initialisation) using the preconditioned loss (12) equipped with 4 different norms: the $L^{1}$ and $L^{2}$ norm of the discrete Riesz map of the residual and its gradient; we denote the $L^{1}$ and $L^{2}$ norms on the gradient as $W^{1,1}$ and $W^{1,2}$ norms, since $W^{k,p}$ represents the Sobolev norm with $k$ derivatives and $p$-th power. In Fig. 4, we report the $L^{2}$ and $H^{1}$ errors as a function of training iteration for the four norms. For reference, we also include the results for the unpreconditioned case, labelled as “N/A”. The first thing to note is that all the preconditioned curves are much steeper than the unpreconditioned ones, indicating that the application of the preconditioner significantly accelerates the convergence of both NNs and their interpolations. Moreover, the $L^{2}$ convergence for both the NN and its interpolation is highly unstable when a preconditioner is not applied, as evidenced by the error curves’ fluctuating pattern. In contrast, the preconditioned error curves are all steadily decreasing, suggesting that the preconditioner effectively stabilises the training process. This also reinforces the conclusion from our previous work [17] that using the $L^{2}$ norm in the preconditioned loss can accelerate the convergence of training in terms of $L^{2}$ error. Furthermore, the $H^{1}$ convergence of the interpolated NNs rapidly decreases in both unpreconditioned and preconditioned cases, as indicated by the stair-like pattern of the convergence curves. However, upon comparing Fig. 4c and Fig. 4d, it becomes apparent that the convergence of NN and its interpolation are not at the same pace. Additional iterations are necessary to ensure the convergence of the NN, even if its interpolation has already reached a favourable position. To sum up, when using the preconditioner $\mathbf{B}_{\rm inv\_lin}$, all the four norms in (12) are effective, with the $W^{1,1}$ norm being slightly better in terms of convergence speed and stability.
@@ 5.2. The 2D Fichera problem with singularity
expressed in polar coordinates. The true solution and the domain are displayed in Fig. 6a. Note that derivatives of $u$ are singular at the origin, in particular, $u\in H^{5/3-\epsilon}(\Omega)$ for any $\epsilon>0$. We have tackled the same problem on the unit square in [17], noting that while the interpolated NNs can recover the FEM solution on a fixed mesh, there is still room for their generalisation to be improved. Therefore, our goal in this experiment is to assess whether the proposed adaptive training strategy improves the NNs’ generalisation.
@@ 5.2.1. The comparison of error indicators
Let us first examine the effectiveness of all three error indicators. Starting with an initial mesh shown in Fig. 6b (resulting from the application of 4 levels of uniform refinement to the initial coarse mesh $\mathcal{C}$), we perform 20 independent runs for each indicator, and present the errors versus the adaptation step in Fig. 7. We choose $k_{U}=4$ in these experiments. The first observation from the figure is that the performance of all three error indicators is comparable, as the lines representing different indicators closely overlap. Besides, for interpolated NNs, the distinction among different error indicators is visually indistinguishable. The $L^{2}$ error curves (Fig. 7a) and $H^{1}$ error curves (Fig. 7c) for all indicators closely follow the red dashed $h$-adaptive FEM line. The variance in errors at each adaptation step is also very low, as indicated by very narrow bands. Additionally, the NNs exhibit strong generalisation in terms of $L^{2}$ error, with all lines in Fig. 7c closely aligning with the FEM line and featuring very narrow bands. More importantly, the $h$-adaptive FEINN method enables NNs to gradually learn the singularity and achieve the $h$-adaptive FEM rate of convergence, as shown in Fig. 7d. Furthermore, the variance in the $H^{1}$ errors for NNs is minimal, as suggested by the narrow bands. In brief, compared to the FEINN method defined on a fixed mesh in our previous work [17], the $h$-adaptive FEINNs method improves the performance of the NNs themselves and enable them to effectively capture the singularity, irrespective of the error indicator employed.
@@ 6. Conclusions
In this work, we propose the $h$-adaptive FEINN method, an extension of the FEINN method [17] that incorporates adaptive meshes. The method inherits the desired properties of FEINNs, namely the applicability to general geometries, simple imposition of Dirichlet boundary conditions and accurate integration of the interpolated weak residual. Furthermore, since the interpolation of the NN is performed on an adaptive FE space, the method preserves the non-linear capabilities of the NN to capture sharp gradients and singularities. The dual norm of the finite element residual in the loss function enhances training robustness and accelerates convergence. We propose a train, estimate, mark and adapt strategy to train the FEINN and adapt the mesh simultaneously. We have carried out a detailed numerical analysis of the method, proving a priori error estimates depending on the expressiveness of the neural network compared to the interpolation mesh.
@@
There are several promising directions for future exploration. One research avenue is to examine the effectiveness of the adaptive training strategy in solving inverse problems. In addition, we aim to investigate the applicability of the proposed approach in attacking problems featuring singularities defined in $H(\mathrm{curl})$ and $H(\mathrm{div})$ spaces [51]. Towards increasing applicability, we would also like to explore the combination of $h$-adaptive FEINNs with unfitted FE spaces [52], which can be used to tackle problems with complex interfaces and geometries. Future work will also focus on extending the framework to $p$ and $r$-refinement strategies, as well as to transient PDEs. Finally, a mathematical analysis of generalisation error bounds for the FEINN-trained NNs is a promising direction for future research.

### v2

@@ 5.1.2. The effect of preconditioning
We consider the preconditioner $\mathbf{B}_{\rm inv\_lin}$, which is defined as the discrete Laplacian resulting from discretisation with the linearised space $V_{h}$ used both as trial and test FE spaces. Thus, $\mathbf{B}_{\rm inv\_lin}$ is a sparse symmetric positive definite (SPD) matrix with reduced computational cost compared to a regular sparse matrix. Note that we apply $k_{U}$ levels of uniform refinement for the partition associated to $U_{h}$ to obtain the mesh for the linearised test space $V_{h}$. Besides, in [Badia2024], we have observed the effectiveness of a GMG preconditioner as an approximation to the inverse of $\mathbf{B}_{\rm inv\_lin}$, suggesting its potential benefits for this problem as well. Nonetheless, for the purpose of this study, we just exactly invert $\mathbf{B}_{\rm inv\_lin}$ (by solving linear systems), as it is sufficient to demonstrate the impact of preconditioning. In these experiments, we cut the number of training iterations by reducing those used in the unpreconditioned experiments by a factor of 4, except for the final adaptation step, where these are divided by a factor of 2. This adjustment accommodates for the increased demand for training on the finest meshes.
@@
In addition to evaluating to what extent preconditioners can expedite convergence, we also investigate the impact of norm choice on the convergence. We train the same NN (i.e., the same parameter initialisation) using the preconditioned loss (12) equipped with 4 different norms: the $L^{1}$ and $L^{2}$ norm of the discrete Riesz map of the residual and its gradient; we denote the $L^{1}$ and $L^{2}$ norms on the gradient as $W^{1,1}$ and $W^{1,2}$ norms, since $W^{k,p}$ represents the Sobolev norm with $k$ derivatives and $p$-th power. In Fig. 4, we report the $L^{2}$ and $H^{1}$ errors as a function of training iteration for the four norms. For reference, we also include the results for the unpreconditioned case, labelled as “N/A”. The first thing to note is that all the preconditioned curves are much steeper than the unpreconditioned ones, indicating that the application of the preconditioner significantly accelerates the convergence of both NNs and their interpolations. Moreover, the $L^{2}$ convergence for both the NN and its interpolation is highly unstable when a preconditioner is not applied, as evidenced by the error curves’ fluctuating pattern. In contrast, the preconditioned error curves are all steadily decreasing, suggesting that the preconditioner effectively stabilises the training process. This also reinforces the conclusion from our previous work [Badia2024] that using the $L^{2}$ norm in the preconditioned loss can accelerate the convergence of training in terms of $L^{2}$ error. Furthermore, the $H^{1}$ convergence of the interpolated NNs rapidly decreases in both unpreconditioned and preconditioned cases, as indicated by the stair-like pattern of the convergence curves. However, upon comparing Fig. 4c and Fig. 4d, it becomes apparent that the convergence of NN and its interpolation are not at the same pace. Additional iterations are necessary to ensure the convergence of the NN, even if its interpolation has already reached a favourable position. To sum up, when using the preconditioner $\mathbf{B}_{\rm inv\_lin}$, all the four norms in (12) are effective, with the $W^{1,1}$ norm being slightly better in terms of convergence speed and stability.
@@ 5.2. The 2D Fichera problem with singularity
expressed in polar coordinates. The true solution and the domain are displayed in Fig. 6a. Note that derivatives of $u$ are singular at the origin, in particular, $u\in H^{5/3-\epsilon}(\Omega)$ for any $\epsilon>0$. We have tackled the same problem on the unit square in [Badia2024], noting that while the interpolated NNs can recover the FEM solution on a fixed mesh, there is still room for their generalisation to be improved. Therefore, our goal in this experiment is to assess whether the proposed adaptive training strategy improves the NNs’ generalisation.
@@ 5.2.1. The comparison of error indicators
Let us first examine the effectiveness of all three error indicators. Starting with an initial mesh shown in Fig. 6b (resulting from the application of 4 levels of uniform refinement to the initial coarse mesh $\mathcal{C}$), we perform 20 independent runs for each indicator, and present the errors versus the adaptation step in Fig. 7. We choose $k_{U}=4$ in these experiments. The first observation from the figure is that the performance of all three error indicators is comparable, as the lines representing different indicators closely overlap. Besides, for interpolated NNs, the distinction among different error indicators is visually indistinguishable. The $L^{2}$ error curves (Fig. 7a) and $H^{1}$ error curves (Fig. 7c) for all indicators closely follow the red dashed $h$-adaptive FEM line. The variance in errors at each adaptation step is also very low, as indicated by very narrow bands. Additionally, the NNs exhibit strong generalisation in terms of $L^{2}$ error, with all lines in Fig. 7c closely aligning with the FEM line and featuring very narrow bands. More importantly, the $h$-adaptive FEINN method enables NNs to gradually learn the singularity and achieve the $h$-adaptive FEM rate of convergence, as shown in Fig. 7d. Furthermore, the variance in the $H^{1}$ errors for NNs is minimal, as suggested by the narrow bands. In brief, compared to the FEINN method defined on a fixed mesh in our previous work [Badia2024], the $h$-adaptive FEINNs method improves the performance of the NNs themselves and enable them to effectively capture the singularity, irrespective of the error indicator employed.
@@ 6. Conclusions
In this work, we propose the $h$-adaptive FEINN method, an extension of the FEINN method [Badia2024] that incorporates adaptive meshes. The method inherits the desired properties of FEINNs, namely the applicability to general geometries, simple imposition of Dirichlet boundary conditions and accurate integration of the interpolated weak residual. Furthermore, since the interpolation of the NN is performed on an adaptive FE space, the method preserves the non-linear capabilities of the NN to capture sharp gradients and singularities. The dual norm of the finite element residual in the loss function enhances training robustness and accelerates convergence. We propose a train, estimate, mark and adapt strategy to train the FEINN and adapt the mesh simultaneously. We have carried out a detailed numerical analysis of the method, proving a priori error estimates depending on the expressiveness of the neural network compared to the interpolation mesh.
@@
There are several promising directions for future exploration. One research avenue is to examine the effectiveness of the adaptive training strategy in solving inverse problems. In addition, we aim to investigate the applicability of the proposed approach in attacking problems featuring singularities defined in $H(\mathrm{curl})$ and $H(\mathrm{div})$ spaces [OLM2019]. Towards increasing applicability, we would also like to explore the combination of $h$-adaptive FEINNs with unfitted FE spaces [Badia2021], which can be used to tackle problems with complex interfaces and geometries. Future work will also focus on extending the framework to $p$ and $r$-refinement strategies, as well as to transient PDEs. Finally, a mathematical analysis of generalisation error bounds for the FEINN-trained NNs is a promising direction for future research.